OpenStudy (sammietaygreen):

Find cos(x) if sin^2 x-1/cos x = -1

OpenStudy (anonymous):

do you know that \[1-\sin^2(x)=\cos^2(x)\]?

OpenStudy (sammietaygreen):

no I am not sure

OpenStudy (anonymous):

oh, well it does

OpenStudy (sammietaygreen):

Oh I thought I had to solve for something. That is just a trig function sorry lol

OpenStudy (anonymous):

the mother of all trig identities is \[\cos^2(x)+\sin^2(x)=1\] and all its cousins for example if you subtract \(\sin^2(x)\) you get \[1-\sin^2(x)=\cos^2(x)\]

OpenStudy (anonymous):

you don't have \[1-\sin^2(x)\]you have \[\sin^2(x)-1\] which is the same as \[-\cos^2(x)\]

OpenStudy (sammietaygreen):

So you basically just have to simplify the trig function but not to a number?

OpenStudy (anonymous):

you still have to solve the equation for \(x\)

OpenStudy (anonymous):

but since the numerator is \(-\cos^2(x)\) it should be relatively easy

OpenStudy (sammietaygreen):

How do I solve for x? I have an idea but it's not like a normal equation

OpenStudy (anonymous):

put \(-\cos^2(x)\) in the top, then cancel a cosine

OpenStudy (sammietaygreen):

on the top of what? like a fraction?

OpenStudy (sammietaygreen):

or set it equal to -1?

OpenStudy (sammietaygreen):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The original equation is this \[\Large \frac{\sin^2(x)-1}{\cos(x)} = -1\] right? or am I off?

jimthompson5910 (jim_thompson5910):

OR is it this \[\Large \sin^2(x)-\frac{1}{\cos(x)} = -1\] @sammietaygreen

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

so as @satellite73 is saying, you'd follow these steps \[\Large \frac{\sin^2(x)-1}{\cos(x)} = -1\] \[\Large \frac{-\cos^2(x)}{\cos(x)} = -1\] \[\Large -\cos(x) = -1\] \[\Large \cos(x) = 1\]

OpenStudy (sammietaygreen):

so do I just solve x when it is simplified to cos(x) = 1

jimthompson5910 (jim_thompson5910):

no need to solve for x itself they just want the value of `cos(x)`

OpenStudy (sammietaygreen):

Ohhh I get it now I don't know why I didn't understand before. So you just had to simplify it to get 1

jimthompson5910 (jim_thompson5910):

yep 1 goes in the box

OpenStudy (sammietaygreen):

Okay. I just didn't understand what he was understanding about x but I get what you're saying. Thank you very much.